# signature_mor.bib

### v1

@@ 2.1 Signatures of smooth paths
where $e_{i}$, $i=1,\ldots d$, denote the standard basis vectors of $V$ as lifted to elements of $T((V))$. The truncated signature satisfies the same ODE, but interpreted on the truncated tensor algebra $T^{m}(V)$, i.e., with a nilpotent tensor product, which will – abusing notation – still be denoted $\otimes$. That is, the truncated signature satisfies
(2.6) $\mathrm{d}\mathbbm{x}_{0,t}^{\leq m}=\mathbbm{x}_{0,t}^{\leq m}\otimes\mathrm{d}x_{t}=\sum_{i=1}^{d}\mathbbm{x}_{0,t}^{\leq m}\otimes e_{i}\mathrm{d}x^{i}_{t},\quad\mathbbm{x}_{0,0}^{\leq m}=1\in T^{m}(V).$
Remark 2.1.
We do not endow $T((V))$ with a topology, and, hence, the ODE (2.5) is only defined in a formal way. Several topologies for $T((V))$ have been considered in the rough path literature, including natural Hilbert or Banach sub-spaces of $T((V))$, see [chevyrev2018signature] for more details. In this paper, we will only ever use the full signature defined on $T((V))$ for motivation, the mathematical analysis will take place on $T^{m}(V)$ for finite truncation $m$. As a finite dimensional vector space, $T^{m}(V)$ will be equipped with the standard Euclidean metric.
The signature is invariant under re-parameterization of the path: I.e., if we consider a (smooth, increasing) map, say, $\gamma:[R,S]\to[0,T]$, and a path $\bar{x}:[R,S]\to V$, $u\mapsto x_{\gamma(u)}$, then $\mathbbm{x}_{\gamma(s),\gamma(t)}^{<\infty}=\bar{\mathbbm{x}}_{s,t}^{<\infty}$, $R\leq s\leq t\leq S$. Up to re-parameterization and so-called tree-like excursions, the signature $\mathbbm{x}^{<\infty}_{0,T}$ uniquely determines the path $x$, see [hambly2010uniqueness]. Invariance under parameterization as well as the possibility of tree-like excursions can be avoided by adding the component $t$ to the path. We will often prefer to work with time-extended paths – i.e., paths with a component $x^{1}_{t}\equiv t$.
(2.7) $\left\langle\ell\,,\mathfrak{a}\right\rangle\coloneqq\sum_{j=1}^{k}\gamma_{j}\left\langle\mathbf{w}_{j}\,,\mathfrak{a}\right\rangle\coloneqq\sum_{j=1}^{k}\gamma_{j}\alpha_{(i^{j}_{1},\ldots,i^{j}_{K_{j}})}\in\mathbb{R}.$
2.2 Signatures for Hölder continuous paths
$\sup_{s<t}\frac{\|x_{t}-x_{s}\|}{|t-s|^{\alpha}}$
is finite. Then, we write $x\in C^{\alpha}$. We briefly sketch the concepts of rough paths and (truncated) signatures for $x\in C^{\alpha}$ – see [friz2020course] for more details. Formally, a rough path $\mathbf{x}$ is a two-parameter function from the simplex $\set{0\leq s\leq t\leq T}$ taking values in the truncated tensor algebra $T^{m}(V)$.
Remark 2.2.
$\mathrm{d}y_{t}=U(y_{t})\mathrm{d}x_{t}.$
However, a formal Taylor expansion shows that higher order Euler approximations of the differential equation in terms of the iterated integral of $x$ of order up to $m=\left\lfloor 1/\alpha\right\rfloor$ would converge. Of course, the catch is that – once again – $C^{\alpha}$ paths are not regular enough for the iterated integrals to make sense in a classical way when $\alpha\leq 1/2$. On the flip side, if we enhance the path increments $x_{s,t}$ with higher order terms $\mathbbm{x}^{I}_{s,t}$ “behaving like iterated integrals” (see [friz2020course] for details), then we can solve the corresponding controlled differential equations for smooth enough vector fields in a pathwise, deterministic way.
$\varrho_{\alpha}(\mathbf{x},\widetilde{\mathbf{x}}):=\sum_{n=1}^{m}\sup_{s<t}\frac{\|\mathbbm{x}^{(n)}_{s,t}-\tilde{\mathbbm{x}}^{(n)}_{s,t}\|}{|t-s|^{n\alpha}},$
where $\mathbf{x}_{s,t}\coloneqq 1+\sum_{n=1}^{m}\mathbbm{x}^{(n)}_{s,t}\in T^{m}(V)$ and $\tilde{\mathbf{x}}$ is defined accordingly – with $\mathbbm{x}^{(1)}_{s,t}=x_{s,t}\in V$ and $\mathbbm{x}^{(n)}_{s,t}\in V^{\otimes n}$, $n=2,\ldots,m$. Now, we can define the desired concepts.
Definition 2.3.
$\mathbf{x}_{s,t}=1+\sum_{n=1}^{N}\mathbbm{x}^{(n)}_{s,t}\in T^{N}(V)$
We denote the set of all geometric $\alpha$-Hölder rough paths by $\mathscr{C}^{\alpha}_{g}([0,T];V)$. We also introduce $\widehat{\mathscr{C}}^{\alpha}_{g}([0,T];V)\coloneqq\Set{\mathbf{x}\in\mathscr{C}^{\alpha}_{g}([0,T];V)}{\forall t\in[0,T]:\,x^{1}_{0,t}=t}$, where $x_{s,t}$ denotes the level-$1$ component of $\mathbf{x}_{s,t}$.
For general $m$, the (truncated) signature $\mathbbm{x}^{\leq m}$ and $\mathbbm{x}^{<\infty}$ of a geometric $\alpha$-Hölder rough path $\mathbf{x}$ can be defined as the limit of $(\mathbbm{x}^{\epsilon})^{\leq m}$ and $(\mathbbm{x}^{\epsilon})^{<\infty}$, respectively, as $\epsilon\to 0$.
Definition 2.4.
(2.8) $\mathrm{d}y(t)=U(y(t))\,\mathrm{d}\mathbf{x}_{t},\quad y(0)=y_{0}\in W,$
exist and converge in the $\alpha$-Hölder metric to $y$.
(2.9) $\mathrm{d}\mathbbm{x}_{0,t}^{\leq m}=\mathbbm{x}_{0,t}^{\leq m}\otimes\mathrm{d}\mathbf{x}_{t}=\sum_{i=1}^{d}\mathbbm{x}_{0,t}^{\leq m}\otimes e_{i}\mathrm{d}\mathbf{x}^{i}_{t},\quad\mathbbm{x}_{0,0}^{\leq m}=1\in T^{m}(V).$
Next, we formulate a result on existence and uniqueness of a solution of (2.8) as well as on properties of the solution map.
Theorem 2.5.
Given $\mathbf{x}\in\mathscr{C}^{\alpha}([0,T],V)$ for $\alpha\in\left(0,\frac{1}{2}\right]$ and $U\in\mathcal{C}_{b}^{N+1}(W,L(V,W))$ or linear, where $N\coloneqq\left\lfloor 1/\alpha\right\rfloor$. Then, there is a unique solution $y\in C^{\alpha}([0,T],W)$ for (2.8). Moreover, the solution map $\mathbf{x}\mapsto g(\mathbf{x})=y$ of (2.8) is locally Lipschitz continuous.
Proof.
We refer to [friz2020course] for the case $\alpha>1/3$. The more general ($p$-variation) framework can be found in [friz2010multidimensional]. ∎
Remark 2.6.
  1. 1.
    the lift of $x$ to a rough path $\mathbf{x}$ (discontinuous);
  2. 2.
    the solution map $\mathbf{x}\mapsto y$ (locally Lipschitz continuous by Theorem 2.5).
This fact is sometimes called the “rough path principle”.
2.3 Signatures of semi-martingales
If, instead of a single deterministic path $x$, we are given a continuous semi-martingale $X:\Omega\times[0,T]\to V$, then the statements above remain true, mutatis mutandis, provided that iterated integrals are defined as Stratonovich integrals rather than Itô integrals.
We define the (truncated) signatures $\mathbb{X}^{<\infty}_{s,t}$, $\mathbb{X}^{\leq m}_{s,t}$ as in (2.2) and (2.4), but with the iterated integrals (2.1) replaced by
(2.10) $\mathbb{X}^{I}_{s,t}\coloneqq\int_{s<t_{1}<\cdots<t_{n}<t}\circ\mathrm{d}X^{i_{1}}_{t_{1}}\otimes\cdots\otimes\circ\mathrm{d}X^{i_{n}}_{t_{n}}.$
Indeed, we can see that the thus defined rough path lift $\mathbf{X}$ a.s. takes values in $\mathscr{C}^{\alpha}_{g}([0,T];V)$ for any $21/3<\alpha<1/$ 2 – we again refer to [friz2020course] for details. Moreover, the solution of a stochastic differential equation, say
$\mathrm{d}Y_{t}=U(Y_{t})\circ\mathrm{d}X_{t},\quad Y_{0}=y_{0}\in W,$
$\mathrm{d}Y_{t}=U(Y_{t})\mathrm{d}\mathbf{X}_{t},\quad Y_{0}=y_{0}\in W,$
where $U:W\to L(V,W)$ is sufficiently smooth.
Remark 2.7.
Remark 2.8.
The setting can also incorporate semimartingales with jumps, see, for instance, [cuchiero2022universal] for a precise statement and proof of universality of signatures in that case.
2.4 Universal approximation theorem
We will now formulate universality of signatures, a well known result in the literature (see, for instance, [KLPA20, bayer2023primal]). Recall that $\widehat{\mathscr{C}}^{\alpha}_{g}([0,T];V)$ denotes the set of geometric $\alpha$-Hölder rough paths $\mathbf{x}$ such that the first component $x^{1}$ of the underlying path $x$ is equal to running time, see Definition 2.3. As noted in the smooth case in Section 2.1, the signature $\mathbbm{x}^{<\infty}_{0,T}$ of a time-extended rough path $\mathbf{x}\in\widehat{\mathscr{C}}^{\alpha}_{g}([0,T];V)$ characterizes $\mathbf{x}$ – and, hence, $x$ up to the initial value $x_{0}$. This, together with the fact that linear functionals of the signature form an algebra – also already mentioned for smooth paths in Section 2.1, but equally true for rough paths – allow us to apply the Stone-Weierstrass theorem.
Theorem 2.9.
Proof of Theorem 2.9..
$\mathcal{A}\coloneqq\Set{\mathbf{x}\mapsto\left\langle\ell\,,\mathbbm{x}^{<\infty}_{0,T}\right\rangle}{\ell\in\mathcal{W}_{d},\ \mathbf{x}\in\widehat{\mathscr{C}}^{\alpha}_{g}([0,T];V)}$
is a subalgebra of $C\left(\widehat{\mathscr{C}}^{\alpha}_{g}([0,T];V);\mathbb{R}\right)$ which is point-separating and contains a non-zero constant function. First note that for any $\ell\in\mathcal{W}_{d}$, the map $\mathbf{x}\mapsto\left\langle\ell\,,\mathbbm{x}^{<\infty}_{0,T}\right\rangle$ is continuous, see, for instance, [friz2010multidimensional]. $\mathcal{A}$ is point-separating since the signature $\mathbbm{x}^{<\infty}$ uniquely determines the rough path $\mathbf{x}$ by the discussion above. The constant function $\mathbf{x}\mapsto 1$ is obviously contained in $\mathcal{A}$ by simply choosing $\ell=\emptyset$, where $\emptyset$ denotes the empty word.
This leaves us to prove that $\mathcal{A}$ is an algebra. In addition to the concatenation product, $\mathcal{W}_{d}$ is also equipped with a commutative shuffle product $\shuffle:\mathcal{W}_{d}\times\mathcal{W}_{d}\to\mathcal{W}_{d}$, see, e.g., [bayer2021optimal], and the following shuffle identity holds for signatures: For $\ell_{1},\ell_{2}\in\mathcal{W}_{d}$ and a geometric rough path $\mathbf{x}$, we have
(2.11) $\left\langle\ell_{1}\,,\mathbbm{x}_{s,t}^{<\infty}\right\rangle\left\langle\ell_{2}\,,\mathbbm{x}_{s,t}^{<\infty}\right\rangle=\left\langle\ell_{1}\shuffle\ell_{2}\,,\mathbbm{x}_{s,t}^{<\infty}\right\rangle.$
Hence, $\mathcal{A}$ is an algebra and the proof is complete. ∎
Example 2.10.
Let $g$ be the solution map of (2.8), i.e., $y=g(\mathbf{x})$, mapping a rough path $\mathbf{x}$ to a path $y$. Then, we know by Theorem 2.5 that this map is continuous. For that reason, a potential continuous functional of interest in Theorem 2.9 can be the $i$th component $y_{i}$ of $y$ evaluated at $T$ (e.g., think of $W=\mathbb{R}^{d}$) meaning that $f(\mathbf{x})=y_{i}(T)$.
As noted in Section 2.3, we can also lift semi-martingales $X$ to the associated rough path $\mathbf{X}$ taking values in $\widehat{\mathscr{C}}^{\alpha}_{g}([0,T];V)$– assuming $X^{1}_{t}\equiv t$ –, and Theorem 2.9 applies:
Corollary 2.11.
We also refer to [cuchiero2022universal, Corollary 3.8] for a more general version applicable to jump-semimartingales.
2.5 Dynamic approximations with signatures
Theorem 2.9 shows that a function $f(x|_{[0,T]})$ of a path [^3] For simplicity, we will consider a smooth path $x$ for the time being, but the discussion is equally valid for the rough case. defined on $[0,T]$ can be approximated by linear functionals of the path’s signature $\left\langle\ell\,,\mathbbm{x}^{<\infty}_{0,T}\right\rangle$ on the interval $[0,T]$. Suppose that we are instead given a path-valued functional, i.e., $y(t)=f(t,x|_{[0,t]})$ – think of a stochastic process $y$ adapted to the filtration generated by $x$. Under the required regularity conditions, Theorem 2.9 immediately implies that we can find approximations
To formalise the problem, we introduce the space of stopped rough paths, going back to [dupire2019functional], see also, [KLPA20, bayer2021optimal], for a more didactic presentation we also refer to [bayer_hager_riedel_book24].
(2.12) $\Lambda_{T}^{\alpha}\coloneqq\bigcup_{t\in[0,T]}\widehat{\mathscr{C}}^{\alpha}_{g}([0,t];V).$
(We restrict ourselves to the time-extended case from the beginning because any such universal approximation result will anyway require us to add time as a component.) Using a similar definition (see [bayer2023primal]), we can define a rough path metric $d$ on $\Lambda^{\alpha}_{T}$, under which it is a Polish space. In this framework, repeating the arguments used in the proof of Theorem 2.9 gives (see [bayer_hager_riedel_book24] for more details)
Corollary 2.12.
Remark 2.13.
The above universal approximation theorems – as typical in the area – are formulated in terms of uniform convergence on compacts. We admit that this concept is problematic, especially in infinite dimensions, as compact sets are often very small. Note that when we consider stochastic processes, e.g., the lift $\mathbf{X}|_{[0,t]}$ of a semimartingale taking values in $\Lambda^{\alpha}_{T}$ (a Polish space), tightness implies that for every $\delta>0$ we can find a compact subset $\mathcal{[}K]\subset\Lambda^{\alpha}_{T}$ such that $P\left(\mathbf{X}\notin\mathcal{K}\right)<\delta$, which allows us to replace “$\epsilon$-close on compact sets” by “$\epsilon$-close with probability $1-\delta$”.
More general global universal approximation results can be formulated in terms of weighted spaces (see [cuchiero2023global]) or in terms of $L^{p}$-norms by using so-called “robust signatures” (see [chevyrev2018signature] for the definition of robust signatures and [bayer2023primal, SA23] for the universality in $L^{p}$.)
2.6 Example: Signature models in finance
As guiding example, we consider signature models for financial markets as introduced in [cuchiero2022signature]. Consider a $d$-dimensional underlying base process $X$, which is assumed to be a (time-extended) continuous semimartingale. Given a linear functional $\ell\in\mathcal{W}_{d}$, we then consider an asset price process given by
(2.13) $S_{t}=S^{(\ell)}_{t}\coloneqq\left\langle\ell\,,\mathbb{X}^{\leq m}_{0,t}\right\rangle,$
where the truncation degree $m$ is chosen high enough that all words in $\ell$ of degree larger than $m$ have zero coefficient. Note that conditions on $\ell$ can be formulated such that the resulting price process satisfies fundamental requirements of mathematical finance such as no-arbitrage. [^4] As such conditions are usually easier to formulate w.r.t. Itô rather than Stratonovich form, those conditions may look simpler in terms of the associated Itô signature. Universality of signatures implies that signature models form a very flexible class of models, capable of approximating many desirable properties of asset price models. In addition, even though the price process (2.13) is not a Markov process, efficient numerical routines for option pricing exist in some cases. [^5] Essentially, when the payoff function can be efficiently approximated by a linear functional of the signature of $S$. As explained in [cuchiero2022signature], the model parameter $\ell$ can be efficiently estimated either from time series of asset prices, or from option price data. However, the dimension of the “state space” $T^{m}(V)$ of (2.13) is generally very large. Indeed, the dimension (in this sense) of the numerical examples in [cuchiero2022signature] are generally of order $100$, but more complex markets might easily lead to much higher dimensional approximations.
In this paper, we assume that we are given a calibrated signature model $S=S^{(\ell)}$ for a fixed linear functional $\ell$ of degree $m$. The purpose is to derive reduced order models, i.e., a linear SDE with solution $\widetilde{\mathbb{X}}$ in $\mathbb{R}^{\widetilde{n}}$ and a linear functional ~$\widetilde{\ell}:\mathbb{R}^{\widetilde{n}}\to\mathbb{R}$ such that
(2.14) ~$\left\langle\widetilde{\ell}\,,\widetilde{\mathbb{X}}\right\rangle\approx S^{(\ell)}\quad\text{and}\quad\widetilde{n}\ll\dim T^{m}(V).$
2.7 Signature differential equation
In the following, we will concentrate on the semi-martingale setting, but note that many discussions are equally valid in the (deterministic or stochastic) rough path framework (or even the bounded variation setting). As seen in (2.6) above, the truncated signature of the semi-martingale $X$ satisfies
@@ 2.7 Signature differential equation
where the tensor product $\otimes$ is truncated to $T^{m}(V)$. Since (2.15) is a linear SDE (driven by a semi-martingale), it fits within the context of stochastic model order reduction, see, for instance, [irka_finance, mor_heston]. For concreteness, we will rewrite the SDE in terms of coordinates. We will use the canonical coordinates
@@
We end up with the SDE
@@
where we use the same symbol $\mathbb{X}^{\leq m}_{0,t}\in\mathbb{R}^{n}$ for the version of the signature in coordinates.
@@ 3 Dimension reduction
As the first step in the analysis, we rewrite (2.16) as an Itô stochastic differential equation. In order to simplify the dimension reduction, we will impose assumptions assuring that the Itô version is still linear.
@@
$X_{t}$ is a time-extended Brownian motion, i.e., $X^{1}_{t}=t$ and $B_{t}\coloneqq(B^{2}_{t},\ldots,B^{d}_{t})^{\top}\coloneqq(X^{2}_{t},\ldots,X^{d}_{t})^{\top}$ is a $(d-1)$-dimensional Brownian motion with associated matrix $K=(k_{ij})_{i,j=2,\dots,d}$ that determines the covariance by $\mathbb{E}[B_{t}B_{t}^{\top}]=Kt$.
@@
Alternatively, we can also consider a general semi-martingale, provided that the quadratic variation processes are also part of $X$. The driving process $X$ is a semi-martingale such that all quadratic co-variations $[X^{i},X^{j}]_{t}$ are already (constant multiples) of coordinates of $X$. More specifically, $X_{t}=(\bar{X}_{t},\widehat{X}_{t})$, where $\bar{X}$ is a (proper) semimartingale and $\widehat{X}$ is a bounded variation process such that for any $i,j$ there is a constant $k_{ij}$ and an index $c_{ij}$ such that $[\bar{X}^{i},\bar{X}^{j}]\equiv k_{ij}\widehat{X}^{c_{ij}}$. Note that Assumption 3.1 is a special case of this semi-martingale setting. For simplicity, we will concentrate on the former case in what follows and leave the situation of this remark for future work.
We aim to rewrite (2.16) in the Itô sense. The associated Itô-Stratonovic correction term is $0.5\sum_{i,j=2}^{d}N_{i}N_{j}k_{ij}$, where $k_{ij}$ is the $ij$-th entry of the matrix $K$ determining the covariance of the Brownian motion. The drift coefficient is $A=N_{1}+0.5\sum_{i,j=2}^{d}N_{i}N_{j}k_{ij}$ leading to
@@
where $z\in\mathbb{R}^{n}$ is a generic initial state and the rows of $L\in\mathbb{R}^{p\times n}$ represent linear functionals of interest. Let us recall that we are particularly interested in $z=\big{(}\begin{matrix}1&0&\dots&0\end{matrix}\big{)}^{\top}$ giving us $\mathcal{X}_{t}=\mathbb{X}^{\leq m}_{0,t}$.

### v2

@@ 2 Signatures
which can be made precise by imposing a suitable topology. In the case of the truncated signature, the corresponding differential equation
(2.2) $\mathrm{d}\mathbbm{x}_{0,t}^{\leq m}=\mathbbm{x}_{0,t}^{\leq m}\otimes\mathrm{d}x_{t}=\sum_{i=1}^{d}\mathbbm{x}_{0,t}^{\leq m}\otimes e_{i}\mathrm{d}x^{i}_{t},\quad\mathbbm{x}_{0,0}^{\leq m}=1\in T^{m}(V),$
is given on the finite-dimensional vector space $T^{m}(V)$, and a proper solution theory is available by, e.g., considering $T^{m}(V)$ as Euclidean space. In both cases, the differential equations should be considered as rough differential equations.
In what follows, we will consider signatures of random paths. More precisely, let $X$ be a continuous semimartingale. Analogously, we define the full and truncated signature $\mathbb{X}^{<\infty}$ and $\mathbb{X}^{\leq m}$ in terms of the Stratonovich integral. We do allow $X$ to have deterministic components, such as $t$ or $\langle X\rangle_{t}$.
In the following, we will concentrate on the semimartingale setting, but note that many discussions are equally valid in the (deterministic or stochastic) rough path framework (or even the bounded variation setting). As seen in (2.2) above, the truncated signature of the semimartingale $X$ satisfies
@@
where the tensor product $\otimes$ is truncated to $T^{m}(V)$. Since (2.3) is a linear SDE (driven by a semimartingale), it fits within the context of stochastic model order reduction, see, for instance, [38, 36]. For concreteness, we will rewrite the SDE in terms of coordinates. We will use the canonical coordinates
@@
Usually, $n$ is very large leading to a high-dimensional linear SDE of the form
@@
where we use the same symbol $\mathbb{X}^{\leq m}_{0,t}\in\mathbb{R}^{n}$ for the version of the signature in coordinates. It is important to notice that $N_{1},\ldots,N_{d}$ do not depend on the underlying semimartingale, but only on the dimension $d$ and the truncation level $m$.
@@
System (2.4) is a large-scale and hence it is computationally expensive to evaluate it. Therefore, we propose an MOR scheme in the following section aiming to reduce computational complexity.
@@ 3 Dimension reduction
We propose a projection-based dimension reduction procedure that can, in principle, be applied directly to the Stratonovich equation (2.4). However, we rewrite (2.4) as an Itô stochastic differential equation for two reasons. First, the Itô case is studied and hence understood much better in the MOR context. Second, there is theoretical evidence that the dimension reduction technique used in this paper performs better when applying it in the Itô framework [27]. In order to simplify the dimension reduction, we will impose assumptions assuring that the Itô version is still linear.
@@
$X_{t}$ is a time-extended Brownian motion, i.e., $X^{1}_{t}=t$ and $B_{t}\coloneqq(B^{2}_{t},\ldots,B^{d}_{t})^{\top}\coloneqq(X^{2}_{t},\ldots,X^{d}_{t})^{\top}$ is a (d1)(d-1)-dimensional Brownian motion with associated matrix $K=(k_{ij})_{i,j=2,\dots,d}$ that determines the covariance by $\mathbb{E}[B_{t}B_{t}^{\top}]=Kt$.
@@
Alternatively, we can also consider a general continuous semimartingale, provided that the quadratic variation processes are also part of $X$. The driving process $X$ is a continuous semimartingale such that all quadratic covariations $X^{i},X^{j}]_{t}$ are already (constant multiples) of coordinates of $X$. More specifically, $X_{t}=(\widehat{X}_{t},\bar{X}_{t})$, where $\bar{X}$ is a (proper) continuous semimartingale and $\widehat{X}$ is a bounded variation process such that for any $i,j$ there is a constant $k_{ij}$ and an index $c_{ij}$ such that $\bar{X}^{i},\bar{X}^{j}]\equiv k_{ij}\widehat{X}^{c_{ij}}$. Note that Assumption 3.1 is a special case of this semimartingale setting. For simplicity, we will concentrate on the former case in what follows and leave the situation of this remark for future work.
We aim to rewrite (2.4) in the Itô sense. The associated Itô-Stratonovich correction term is $0.5\sum_{i,j=2}^{d}N_{i}N_{j}k_{ij}$, where $k_{ij}$ is the $ij$-th entry of the matrix $K$ determining the covariance of the Brownian motion. The drift coefficient is $A=N_{1}+0.5\sum_{i,j=2}^{d}N_{i}N_{j}k_{ij}$ leading to
@@
where $z\in\mathbb{R}^{n}$ is a generic initial state. In particular, we are interested in $z=\big(\begin{matrix}1&0&\dots&0\end{matrix}\big)^{\top}$ giving us the signature process $\mathcal{X}_{t}=\mathbb{X}^{\leq m}_{0,t}$, but we discuss the dimension reduction procedure in more generality. In principle, it is also possible to consider the same MOR scheme yielding an accurate estimate for an entire subspace of initial states but this is beyond the focus of this paper. Moreover, we do not aim to approximate the full state variable $\mathcal{X}_{t}$ when applying a dimension reduction technique. We rather find a reduced model that is able to reproduce a linear map of $\mathcal{X}_{t}$ represented by a matrix $L\in\mathbb{R}^{p\times n}$, where $p$ is potentially much smaller than $n$. This framework is motivated by the universal approximation property of the signature that tells us that nonlinear time dependent dynamics like rough or stochastic differential equations can be approximated by a linear map of the signature, see Sections A.4 and particularly A.5.
Remark 3.3.
Considering the general continuous semimartingale setting, using the notation of Remark 3.2 and assuming that $\widehat{X}$ takes values in $\mathbb{R}^{\hat{d}}$ the Itô SDE in (3.1) is replaced by
$\mathrm{d}\mathcal{X}_{t}=\sum_{k=1}^{\hat{d}}N_{k}\mathcal{X}_{t}\mathrm{d}\widehat{X}_{t}^{k}+0.5\sum_{i,j=\hat{d}+1}^{d}N_{i}N_{j}\mathcal{X}_{t}k_{ij}\mathrm{d}\widehat{X}_{t}^{c_{ij}}+\sum_{k=\hat{d}+1}^{d}N_{k}\mathcal{X}_{t}\mathrm{d}\bar{X}_{t}^{k}.$
Given that $\bar{X}$ is a Brownian motion, we have $\hat{d}=1$ and $\widehat{X}_{t}=t$ yielding (3.1).